# From paintbrush to pixel: A review of deep neural networks in AI-generated art

### v2

@@ 3.3 Text-to-Image Models
Fig. 13: Comparison of the Transformer-based text-to-image models.
@@ 3.3.1 The Transformer
DALL-E. Early in 2021, OpenAI released their groundbreaking model DALL-E (named after Pixar’s Wall-e and Surrealist painter Salvador Dali) on their blog. Shortly after, they detailed the workings of their model in their paper titled “Zero-shot Text-to-Image generation” Ramesh et al. (2021). DALL-E combines a discrete variational autoencoder (dVAE) Rolfe (2017), which learns to map images to lower dimensional tokens, and a Transformer, which autoregressively models text and image tokens (see Fig. 13). In this manner, DALL-E is optimized to jointly model text, accompanying images and their token representations. Given a text description as input, DALL-E can predict the image tokens and decode them into an image during inference. Here, zero-shot refers to the ability to generate images based on text descriptions that are not seen during training. This means that DALL-E can combine concepts that it has learned separately but never seen together in a single generated image. For example, it has seen both robots and illustrations of dragons in the training data but it has not seen a robot in the shape of a dragon. However, when prompted to generate “a robot dragon”, the model can produce sensible images (see Fig. 14). This remarkable capability of the model has resulted in hype surrounding DALL-E. Although DALL-E can generate cartoons and art styles quite well, it lacks accuracy when generating photo-realistic images. As a result, OpenAI and other companies have devoted substantial resources to create an improved text-to-image model.
@@
Fig. 14: Example images produced by DALL-E Ramesh et al. (2021), retrieved from the original OpenAI blog https://openai.com/blog/dall-e/.
@@
CogView. Concurrently with DALL-E, Ding et al. (2021) created CogView, a similar text-to-image model that supports Chinese rather than English. Besides the innovative idea to combine VAE and transformers, they include other features such as super-resolution to improve the resolution of the generated images. Despite their super-resolution module, their generated images lack photo-realistic quality.
@@
Make-A-Scene. In 2022, Meta AI released their Make-A-Scene model Gafni et al. (2022). Their Transformer-based text-to-image model allows the user more control over the generated image by working with segmentation maps. During training, the model receives a text prompt, segmentation map and accompanying image as input (similarly as GauGAN2). The model then learns a latent mapping based on tokenized versions of the inputs. During inference, Make-A-Scene is able to generate an image and segmentation map based solely on text input (see Fig. 13). Alternatively, one can provide a segmentation map as input to steer the desired output. Moreover, one can alter the segmentation map that the model produces to steer the image generation.
@@
Parti. Later in 2022, Google released their Transformer-based text-to-image model called Parti which stands for Pathways Autoregressive Text-to-Image model Yu et al. (2022a). This was the second text-to-image model Google released that year, a month after releasing Imagen Saharia et al. (2022) (see Section 3.3.2). Parti is based on a ViT-VQGAN Yu et al. (2022b), which combines the transformer encoder and decoder with an adversarial loss of a pretrained GAN to optimize image generation. Parti uses an additional transformer encoder to handle the text input, which is transformed into text tokens to serve as input to a transformer decoder alongside the image tokens from the ViT-VQGAN during training. At test time, the Transformer solely receives text as input and predicts the image tokens, which are then provided to the ViT-VQGAN to detokenize and turn them into an actual image (see Fig. 13). Parti outperforms all other text-to-image models on FID score, with a zero-shot FID score of 7.23 and a fine-tuned FID score of 3.22 on MS-COCO Lin et al. (2014) (see Table 1). Fig. 15 is a great example of Parti’s remarkable capability to extract the essence of what the caption refers to (in this case, the style of American Gothic), rather than simply generating a copy of the original in which the rabbits seem photoshopped.
@@
Fig. 15: (Left) American Gothic by Grant Wood. (Right) An example image generated by Parti Yu et al. (2022a), retrieved from https://parti.research.google/.
@@
Muse. Early in 2023, Google released Muse Chang et al. (2023), another Transformer-based text-to-image model. Muse includes a VQGAN tokenizer to transfer images into tokens and vice versa. The text input is turned into tokens using a pre-trained transformer encoder of the T5-XXL language model Raffel et al. (2020) (see Fig. 13). The advantage of using such a pre-trained encoder is that this model component has been trained on a large corpus of text. One struggle in training text-to-image models is the time consuming gathering of a large set of high quality image and caption pairs. Therefore, the text/caption pairs likely only span a limited section of the known concepts within a language. Since NLP models are trained solely with text as input (which is easier to gather), they are trained on a more encompassing corpus of text. The authors find that the inclusion of the pretrained encoder results in higher quality, photo-realistic images and that the generation of text in the generated images is more accurate. Since then, other text-to-image models have incorporated T5-XXL as text encoder in a similar manner (see Section 3.3.2 on diffusion models).
@@
Muse achieves a CLIP score of 0.32, which is a measure of how close the generated images are to the prompted caption. Chang et al. (2023) verified this further in a behavioral study in which human raters indicated that the Muse-generated images are better aligned with the prompts compared to images generated with Stable Diffusion (see Section 3.3.2). In addition to image generation, Muse allows inpainting, outpainting and mask-free editing (which will be explained in more detail in Section 3.3.2).
@@ 3.3.2 Diffusion Models
GLIDE. In 2021, OpenAI came out with a paper that showed that diffusion models outperform GANs on image generation Dhariwal and Nichol (2021). Less than a year after, OpenAI applied this insight to text-to-image generation, and they released GLIDE Nichol et al. (2021), a pipeline consisting of a diffusion model for image synthesis and a transformer encoder for text input (see Fig. 17). This new and improved model is trained on the same dataset as DALL-E, yet the quality of their generated images strongly outperforms DALL-E. In a study where they asked human participants to judge the generated images from DALL-E and GLIDE, the raters preferred the GLIDE images over the DALL-E images 87% of the time for photorealism and 69% of the time for caption similarity. In addition, they preferred blurred GLIDE images over reranked or cherry picked DALL-E images.
@@
Fig. 16: An example of inpainting with GLIDE. It receives an image of the Mona Lisa as input as well as a mask and a text description “a golden necklace”. Then, it generates the output as Mona Lisa with a golden necklace. Nichol et al. (2021).
@@
DALL-E 2. Even though GLIDE was an impressive improvement upon DALL-E, it did not garner the same attention. When OpenAI released DALL-E 2 Ramesh et al. (2022) in 2022, an advancement of GLIDE, this changed. DALL-E 2 has a similar diffusion pipeline as GLIDE, but the text input of this diffusion pipeline has been improved. Whereas GLIDE uses an untrained transformer encoder to format the text, DALL-E 2 uses the CLIP text encoder. Additionally, it uses a prior model to transform the text embedding into a CLIP image embedding before feeding it to the diffusion model (see Fig. 17). This is achieved by transforming images and text descriptions to image and text embeddings (or tokens) respectively using transformer encoders. The model is then trained to link the correct embeddings with one another. The relationship between text descriptions and images is exploited in DALL-E 2, providing the diffusion model with an embedding that reflects the text input while being better suited for image generation. In addition to improving image quality compared to GLIDE, DALL-E 2 allows the user to extend the background of an existing image or computer-generated one (referred to as outpainting, see Fig. 18) and to generate variations of images.
Fig. 17: Comparison of the diffusion model-based text-to-image models.
DALL-E 3. An important limitation of current text-to-image models is their difficulty in accurately reflecting everything in the text prompt they receive. A common example to illustrate their lacking prompt following ability is the generation of a pinguin cartoon, wearing clothing items in different colors. The models tend to have difficulty linking the right color to the right clothing item. Researchers at OpenAI hypothesized that this limitation stems from a lack of correspondence between text descriptions and images in the datasets used to train these text-to-image models. They trained an image captioner to re-label these datasets with more accurate yet synthetic prompts and found that this approach improved the current text-to-image models. DALL-E 3 Betker et al. (2023) incorporates this re-captioning, amongst other model improvements (such as replacing the text encoder with T5-XXL Raffel et al. (2020), similarly as Muse Chang et al. (2023)). At inference time, the model generates an image and a synthetic, rich prompt. The user can then edit the prompt to have more control over the output image. DALL-E 3 is currently available through ChatGPT and Bing image creator.
@@
Fig. 18: Image outpainting examples by DALL-E 2 Ramesh et al. (2022). (Left) Mona Lisa. (Right) Girl with a Pearl Earring.
@@
Imagen. Shortly after DALL-E 2, Google released their first text-to-image model called Imagen Saharia et al. (2022). Their model is closer to GLIDE in architecture, since it does not rely on CLIP embeddings. Instead, it uses the pretrained encoder of the NLP model T5-XXL (similarly as their earlier model Muse), whose embeddings are fed to the diffusion model (see Fig. 17). As a result, the model is able to generate images that contain text more accurately (which is something OpenAI’s models struggled with). In addition, Imagen feeds the generated image of the diffusion model to two super-resolution diffusion models to increase the resolution of the images.
@@
Stable Diffusion. The biggest revolution in the field is perhaps the fully open source release of Stable Diffusion Rombach et al. (2022) by a company called Stability AI (we elaborate on the concerns regarding open source models in Section 5). Their main contribution is the computational efficiency of their model, as opposed to the above mentioned text-to-image models. Rather than operating in pixel-space, Stable Diffusion operates in (lower-dimensional) latent space and maps the output of the diffusion process back to pixel space using a VAE (see Fig. 17). Whereas previous text-to-image models require hundreds of GPU computing days, this latent diffusion model requires significantly smaller computational demands and is therefore more accessible to those with less resources. Besides image generation, Stable Diffusion additionally allows the user to modify existing images through image-to-image translation (e.g., turning a sketch into digital art) or inpainting (removing or adding elements in an existing image).
@@
Model FID on MS-COCO Trainable parameters Open source Capabilities
ArtGAN [Tan et al. (2017)] -* - Yes Image generation
CAN [Elgammal et al. (2017)] - - Yes Image generation
pix2pix [Isola et al. (2017)] - - Yes Image manipulation
CycleGAN [Zhu et al. (2017)] - - Yes Image manipulation
GauGAN [Park et al. (2019)] 22.6 - Yes Image generation, image manipulation, text-to-image
LAFITE [Zhou et al. (2021)] 26.94 75M Yes Text-to-image
DALL-E [Ramesh et al. (2021)] 27.50 12B No Text-to-image
CogView [Ding et al. (2021)] 27.10 4B Yes Text-to-image
Make-A-Scene [Gafni et al. (2022)] 11.84 4B No Text-to-image, image manipulation
Parti [Yu et al. (2022a)] 7.23 20B No Text-to-image
Muse [Chang et al. (2023)] 7.88 3B No Text-to-image, inpainting, outpainting, mask-free editing
GLIDE [Nichol et al. (2021)] 12.24 3.5B Partially Text-to-image, inpainting
DALL-E 2 [Ramesh et al. (2022)] 10.39 4.5B Partially Text-to-image, inpainting, outpainting, generate variations
DALL-E 3 [Betker et al. (2023)] - - Partially Text-to-image, inpainting, outpainting, generate variations
Imagen [Saharia et al. (2022)] 7.27 2B No Text-to-image
Stable Diffusion [Rombach et al. (2022)] 12.63 1.45B Yes Text-to-image, inpainting, image manipulation
SDXL [Podell et al. (2023)] Various 2.6B Yes Text-to-image, inpainting, image manipulation
InstructPix2Pix [Brooks et al. (2023)] - - Yes Mask-free editing
@@
SDXL. Recently, Stability AI released a more powerful and larger version of stable diffusion, called SDXL Podell et al. (2023). This model still operates in latent space but has significantly more parameters than the earlier stable diffusion releases. This is due to the use of more powerful pre-trained text encoders as well as an increase in transformer blocks in the U-Net part of the diffusion model. In addition, SDXL features a new component called the ‘refiner’, which is a U-Net placed in between the output of the latent diffusion and the VAE to map the latent representation to pixelspace. By placing a refiner in between these two steps, the model learns to further refine the image latent before it is mapped back to pixel space. They illustrate that this addition results in images with more fine-grained details, especially in the background. Besides these changes to the diffusion models architecture, they also add new features that the model is conditioned on. This conditioning allows more control over the generated images. For example, SDXL is conditioned on image size and cropping parameters, allowing the user to specify the desired resolution of the output image and how centered the central object should be. As such, SDXL can generate images with a different height and width, which is common in artworks and photographs.
To mitigate the limitation that diffusion models perform multiple steps at inference time to generate an image from noise, various distillation methods have been proposed to generate images with few steps. These distillation methods operated in pixel-space, as most initial diffusion models do. Stability AI proposed a latent distillation method to go with their latent diffusion models that is more computationally efficient Sauer et al. (2024). They coined these models SD Turbo and SDXL Turbo. Their Turbo models are on par or outperform other leading text-to-image models (without distillation method) on image quality and prompt alignment as assessed with a user study.
Fig. 19: InstructPix2Pix turns the Sunflowers by Vincent Van Gogh into a painting of roses Brooks et al. (2023).
@@
Fig. 20: Images generated by DALL-E 3, SDXL, and Midjourney using the same prompt. (Top row) A cat in a detailed Renaissance cityscape with historical architecture. (Middle row) Highly aesthetic versions of the top row. (Bottom row) Less aesthetic versions of the top row, potentially criticized by art critics.
InstructPix2Pix. Although several text-to-image models have the inpainting feature, in practice it is difficult to get the desired output based on a text description. The user either needs to describe the entire output image, or create a mask for the area in the image that should be modified and describe the desired modification. InstructPix2Pix Brooks et al. (2023) is a modification of Stable Diffusion and allows the user to modify images through intuitive text instructions. Rather than having to describe the desired output image or providing a mask, the user can just write an intuitive instruction of how the input image should be adjusted (mask-free editing). For example, if you would like to turn the Sunflowers by Vincent Van Gogh into a painting of roses, you can just write the instruction “Swap sunflowers with roses” (Fig. 19).
4 Comparison of deep generative models
Due to the wide variety of models that have been recently developed in the field of generative AI, it may be hard to decide which model is most suited for one’s purposes. Table 1 provides a comparison of the above mentioned models in terms of their size (trainable parameters), the quality of the generated images (FID score), the capabilities of the model and its accessibility (open source vs. not for public use). It should be noted that FID score Heusel et al. (2017) is a suboptimal measure of image quality and especially image aesthetics. In fact, Kirstain et al. (2023) found that FID score correlates negatively with human preferences for prompt-based generated images. They propose an alternative measure, called PickScore, that correlates highly with human preferences.
@@ 4 Comparison of deep generative models
Fig. 21: Example images generated by SDXL and Midjourney, showcasing the capability to create (partially or fully) nude images. (Left) Images generated by SDXL. (Right) Images generated by Midjourney. The images in the first and second rows were generated with the prompt “A painting of Snow White in the style of The Birth of Venus by Sandro Botticelli.” The images in the third and fourth rows were generated with the prompt “Batman as a Greek God wearing a loincloth.”
To compare the models further on their capability to generate art-like images, we asked the three widely used models to generate images that represent Renaissance paintings. In addition to DALL-E 3 Betker et al. (2023) and SDXL Podell et al. (2023), we include Midjourney in our comparison due to its popularity in text-to-image generation, despite its lack of publicly shared neural network architectures and code. For our comparison, we use prompts to generate images depicting a cat in a cityscape as a Renaissance painting: one highly aesthetic version and another that could potentially be criticized by art critics. Our aim is to see how well these models capture art styles in relation to aesthetics and art criticism. Fig. 20 shows the results. All three models seem to capture the Renaissance style for the cityscape aspect of the image but seem to have difficulty applying this style to the cat. This could be explained by cityscape being a common art genre also present in Renaissance paintings. On the other hand, cats are less common in cityscape or Renaissance paintings but are highly common in natural images as part of these models’ training dataset. This could explain why the cats are often generated in a more modern drawing style and are disproportionately large compared to the buildings in the images. When we compare the highly aesthetic variants to the neutral prompt images, we can see that the highly aesthetic ones are often enhanced in color, brightness or saturation. Music et al. (2023) found that these image features are positively related to human preferences for images generated with a model that aims to enhance the aesthetic value of an image. In a similar line of research, the importance of generating highly aesthetic images has been emphasized by Dai et al. (2023), who propose quality-tuning a latent diffusion model Rombach et al. (2022) for this purpose. In Fig. 20, the images for the art criticism prompts cannot be distinguished easily from the neutral prompts, indicating that the models have a weak notion of art criticism, or that they (most probably correctly) assume that any AI-generated image would be criticized by art critics.

### v3

@@ 3.3 Text-to-Image Models
Figure 13: Comparison of the Transformer-based text-to-image models.
@@ 3.3.1 The Transformer
DALL-E. Early in 2021, OpenAI released their groundbreaking model DALL-E (named after Pixar’s Wall-e and Surrealist painter Salvador Dali) on their blog. Shortly after, they detailed the workings of their model in their paper titled “Zero-shot Text-to-Image generation” [85]. DALL-E combines a discrete variational autoencoder (dVAE) [89], which learns to map images to lower dimensional tokens, and a Transformer, which autoregressively models text and image tokens (see Fig. 13). In this manner, DALL-E is optimized to jointly model text, accompanying images and their token representations. Given a text description as input, DALL-E can predict the image tokens and decode them into an image during inference. Here, zero-shot refers to the ability to generate images based on text descriptions that are not seen during training. This means that DALL-E can combine concepts that it has learned separately but never seen together in a single generated image. For example, it has seen both robots and illustrations of dragons in the training data but it has not seen a robot in the shape of a dragon. However, when prompted to generate “a robot dragon”, the model can produce sensible images (see Fig. 14). This remarkable capability of the model has resulted in hype surrounding DALL-E. Although DALL-E can generate cartoons and art styles quite well, it lacks accuracy when generating photo-realistic images. As a result, OpenAI and other companies have devoted substantial resources to create an improved text-to-image model.
@@
Figure 14: Example images produced by DALL-E [85], retrieved from the original OpenAI blog https://openai.com/blog/dall-e/.
@@
CogView. Concurrently with DALL-E, [90] created CogView, a similar text-to-image model that supports Chinese rather than English. Besides the innovative idea to combine VAE and transformers, they include other features such as super-resolution to improve the resolution of the generated images. Despite their super-resolution module, their generated images lack photo-realistic quality.
@@
Make-A-Scene. In 2022, Meta AI released their Make-A-Scene model [91]. Their Transformer-based text-to-image model allows the user more control over the generated image by working with segmentation maps. During training, the model receives a text prompt, segmentation map and accompanying image as input (similarly as GauGAN2). The model then learns a latent mapping based on tokenized versions of the inputs. During inference, Make-A-Scene is able to generate an image and segmentation map based solely on text input (see Fig. 13). Alternatively, one can provide a segmentation map as input to steer the desired output. Moreover, one can alter the segmentation map that the model produces to steer the image generation.
@@
Parti. Later in 2022, Google released their Transformer-based text-to-image model called Parti which stands for Pathways Autoregressive Text-to-Image model [92]. This was the second text-to-image model Google released that year, a month after releasing Imagen [93] (see Section 3.3.2). Parti is based on a ViT-VQGAN [94], which combines the transformer encoder and decoder with an adversarial loss of a pretrained GAN to optimize image generation. Parti uses an additional transformer encoder to handle the text input, which is transformed into text tokens to serve as input to a transformer decoder alongside the image tokens from the ViT-VQGAN during training. At test time, the Transformer solely receives text as input and predicts the image tokens, which are then provided to the ViT-VQGAN to detokenize and turn them into an actual image (see Fig. 13). Parti outperforms all other text-to-image models on FID score, with a zero-shot FID score of 7.23 and a fine-tuned FID score of 3.22 on MS-COCO [95] (see Table 1). Fig. 15 is a great example of Parti’s remarkable capability to extract the essence of what the caption refers to (in this case, the style of American Gothic), rather than simply generating a copy of the original in which the rabbits seem photoshopped.
@@
Figure 15: (Left) American Gothic by Grant Wood. (Right) An example image generated by Parti [92], retrieved from https://parti.research.google/.
@@
Muse. Early in 2023, Google released Muse [96], another Transformer-based text-to-image model. Muse includes a VQGAN tokenizer to transfer images into tokens and vice versa. The text input is turned into tokens using a pre-trained transformer encoder of the T5-XXL language model [97] (see Fig. 13). The advantage of using such a pre-trained encoder is that this model component has been trained on a large corpus of text. One struggle in training text-to-image models is the time consuming gathering of a large set of high quality image and caption pairs. Therefore, the text/caption pairs likely only span a limited section of the known concepts within a language. Since NLP models are trained solely with text as input (which is easier to gather), they are trained on a more encompassing corpus of text. The authors find that the inclusion of the pretrained encoder results in higher quality, photo-realistic images and that the generation of text in the generated images is more accurate. Since then, other text-to-image models have incorporated T5-XXL as text encoder in a similar manner (see Section 3.3.2 on diffusion models).
@@
Muse achieves an FID score of 7.88 along with a CLIP score of 0.32, which is a measure of how close the generated images are to the prompted caption. Different from the FID score, which measures image realism by comparing the feature distributions of generated and real images, the CLIP score evaluates text–image alignment. It leverages the CLIP model to embed text and images into a shared space and quantifies how well a generated image matches the input prompt, with higher scores indicating stronger semantic consistency. [96] verified this further in a behavioral study in which human raters indicated that the Muse-generated images are better aligned with the prompts compared to images generated with Stable Diffusion (see Section 3.3.2). In addition to image generation, Muse allows inpainting, outpainting and mask-free editing (which will be explained in more detail in Section 3.3.2).
@@ 3.3.2 Diffusion Models
GLIDE. In 2021, OpenAI came out with a paper that showed that diffusion models outperform GANs on image generation [58]. Less than a year after, OpenAI applied this insight to text-to-image generation, and they released GLIDE [98], a pipeline consisting of a diffusion model for image synthesis and a transformer encoder for text input (see Fig. 18). This new and improved model is trained on the same dataset as DALL-E, yet the quality of their generated images strongly outperforms DALL-E. In a study where they asked human participants to judge the generated images from DALL-E and GLIDE, the raters preferred the GLIDE images over the DALL-E images 87% of the time for photorealism and 69% of the time for caption similarity. In addition, they preferred blurred GLIDE images over reranked or cherry picked DALL-E images.
@@
Figure 16: An example of inpainting with GLIDE. It receives an image of the Mona Lisa as input as well as a mask and a text description “a golden necklace”. Then, it generates the output as Mona Lisa with a golden necklace. [98].
@@
DALL-E 2. Even though GLIDE was an impressive improvement upon DALL-E, it did not garner the same attention. When OpenAI released DALL-E 2 [99] in 2022, an advancement of GLIDE, this changed. DALL-E 2 has a similar diffusion pipeline as GLIDE, but the text input of this diffusion pipeline has been improved. Whereas GLIDE uses an untrained transformer encoder to format the text, DALL-E 2 uses the CLIP text encoder. Additionally, it uses a prior model to transform the text embedding into a CLIP image embedding before feeding it to the diffusion model (see Fig. 18). This is achieved by transforming images and text descriptions to image and text embeddings (or tokens) respectively using transformer encoders. The model is then trained to link the correct embeddings with one another. The relationship between text descriptions and images is exploited in DALL-E 2, providing the diffusion model with an embedding that reflects the text input while being better suited for image generation. In addition to improving image quality compared to GLIDE, DALL-E 2 allows the user to extend the background of an existing image or computer-generated one (referred to as outpainting, see Fig. 17) and to generate variations of images.
@@
Figure 17: Image outpainting examples by DALL-E 2 [99]. (Left) Mona Lisa. (Right) Girl with a Pearl Earring.
Figure 18: Comparison of the diffusion model-based text-to-image models.
DALL-E 3. An important limitation of current text-to-image models is their difficulty in accurately reflecting everything in the text prompt they receive. A common example to illustrate their lacking prompt following ability is the generation of a penguin cartoon, wearing clothing items in different colors. The models tend to have difficulty linking the right color to the right clothing item. Researchers at OpenAI hypothesized that this limitation stems from a lack of correspondence between text descriptions and images in the datasets used to train these text-to-image models. They trained an image captioner to re-label these datasets with more accurate yet synthetic prompts and found that this approach improved the current text-to-image models. DALL-E 3 [8] incorporates this re-captioning, amongst other model improvements (such as replacing the text encoder with T5-XXL [97], similarly as Muse [96]). At inference time, the model generates an image and a synthetic, rich prompt. The user can then edit the prompt to have more control over the output image. DALL-E 3 is currently available through ChatGPT and Bing image creator.
@@
Imagen. Shortly after DALL-E 2, Google released their first text-to-image model called Imagen [93]. Their model is closer to GLIDE in architecture, since it does not rely on CLIP embeddings. Instead, it uses the pretrained encoder of the NLP model T5-XXL (similarly as their earlier model Muse), whose embeddings are fed to the diffusion model (see Fig. 18). As a result, the model is able to generate images that contain text more accurately (which is something OpenAI’s models struggled with). In addition, Imagen feeds the generated image of the diffusion model to two super-resolution diffusion models to increase the resolution of the images.
@@
Stable Diffusion. The biggest revolution in the field is perhaps the fully open source release of Stable Diffusion [100] by a company called Stability AI (we elaborate on the concerns regarding open source models in Section 5). Their main contribution is the computational efficiency of their model, as opposed to the above mentioned text-to-image models. Rather than operating in pixel-space, Stable Diffusion operates in (lower-dimensional) latent space and maps the output of the diffusion process back to pixel space using a VAE (see Fig. 18). Whereas previous text-to-image models require hundreds of GPU computing days, this latent diffusion model requires significantly smaller computational demands and is therefore more accessible to those with less resources. Besides image generation, Stable Diffusion additionally allows the user to modify existing images through image-to-image translation (e.g., turning a sketch into digital art) or inpainting (removing or adding elements in an existing image).
SDXL. Recently, Stability AI released a more powerful and larger version of Stable Diffusion, called SDXL [101]. This model still operates in latent space but has significantly more parameters than the earlier Stable Diffusion releases. This is due to the use of more powerful pre-trained text encoders as well as an increase in transformer blocks in the U-Net part of the diffusion model. In addition, SDXL features a new component called the ‘refiner’, which is a U-Net placed in between the output of the latent diffusion and the VAE to map the latent representation to pixelspace. By placing a refiner in between these two steps, the model learns to further refine the image latent before it is mapped back to pixel space. They illustrate that this addition results in images with more fine-grained details, especially in the background. Besides these changes to the diffusion models architecture, they also add new features that the model is conditioned on. This conditioning allows more control over the generated images. For example, SDXL is conditioned on image size and cropping parameters, allowing the user to specify the desired resolution of the output image and how centered the central object should be. As such, SDXL can generate images with a different height and width, which is common in artworks and photographs.
To mitigate the limitation that diffusion models perform multiple steps at inference time to generate an image from noise, various distillation methods have been proposed to generate images with few steps. These distillation methods operated in pixel-space, as most initial diffusion models do. Stability AI proposed a latent distillation method to go with their latent diffusion models that is more computationally efficient [102]. They coined these models SD Turbo and SDXL Turbo. Their Turbo models are on par or outperform other leading text-to-image models (without distillation method) on image quality and prompt alignment as assessed with a user study.
FLUX. It is a text-to-image model developed by Black Forest Labs, a start-up based in Germany that was founded by researchers previously involved in the development of various Stable Diffusion versions. FLUX is best known as the image generation engine behind Grok, X’s (formerly Twitter) built-in image generator. The model is a rectified flow transformer, which operates similarly to a diffusion model but with greater efficiency [103]. Traditional diffusion models typically require many iterative steps to transform noise into a data sample. In contrast, rectified flow models reduce the number of required steps by learning a straight line between noise to data. Traditional diffusion models learn by gradually adding Gaussian noise to an image and then learning to denoise it. The Gaussian noise is stochastic, resulting in a curved path between noise and data that requires many steps. On the other hand, rectified flow transformers construct a synthetic path between noise and data. Their learning objective is to transform random noise into data by traveling along these straight line trajectories and learning a velocity vector. Each step involves a linear interpolation between noise and data, enabling faster inference with fewer steps. FLUX offers inpainting, outpainting, and image variation in addition to text-to-image generation. It outperforms previous models in quantitative analyses and human preference ratings.
InstructPix2Pix. Although several text-to-image models have the inpainting feature, in practice it is difficult to get the desired output based on a text description. The user either needs to describe the entire output image, or create a mask for the area in the image that should be modified and describe the desired modification. InstructPix2Pix [104] is a modification of Stable Diffusion and allows the user to modify images through intuitive text instructions. Rather than having to describe the desired output image or providing a mask, the user can just write an intuitive instruction of how the input image should be adjusted (mask-free editing). For example, if you would like to turn the Sunflowers by Vincent Van Gogh into a painting of roses, you can just write the instruction “Swap sunflowers with roses” (Fig. 19).
Figure 19: InstructPix2Pix turns the Sunflowers by Vincent Van Gogh into a painting of roses [104].
4 Comparison of deep generative models
Due to the wide variety of models that have been recently developed in the field of generative AI, it may be hard to decide which model is most suited for one’s purposes. Table 1 provides a comparison of the above mentioned models in terms of their size (trainable parameters), the quality of the generated images (FID score), the capabilities of the model and its accessibility (open source vs. not for public use). It should be noted that FID score [105] is a suboptimal measure of image quality and especially image aesthetics. In fact, [106] found that FID score correlates negatively with human preferences for prompt-based generated images. They propose an alternative measure, called PickScore, that correlates highly with human preferences. We should also note that large diffusion models require substantial computational resources and energy, raising sustainability concern [107].
@@ 4 Comparison of deep generative models
Model
FID on
MS-COCO
Trainable
parameters
Open
source
Capabilities
ArtGAN [80] -* - Yes
Image generation
CAN [81] - - Yes
Image generation
pix2pix [40] - - Yes
Image manipulation
CycleGAN [41] - - Yes
Image manipulation
GauGAN [83] 22.6 - Yes
Image generation,
image manipulation,
text-to-image
LAFITE [86] 26.94 75M Yes
Text-to-image
DALL-E [85] 27.50 12B No
Text-to-image
CogView [90] 27.10 4B Yes
Text-to-image
Make-A-Scene [91] 11.84 4B No
Text-to-image,
image manipulation
Parti [92] 7.23 20B No
Text-to-image
Muse [96] 7.88 3B No
Text-to-image, inpainting,
outpainting, mask-free editing
GLIDE [98] 12.24 3.5B Partially
Text-to-image, inpainting
DALL-E 2 [99] 10.39 4.5B Partially
Text-to-image, inpainting,
outpainting, generate variations
DALL-E 3 [8] - - Partially
Text-to-image, inpainting,
outpainting, generate variations
Imagen [93] 7.27 2B No
Text-to-image
Stable Diffusion [100] 12.63 1.45B Yes
Text-to-image, inpainting,
image manipulation
SDXL [101] Various 2.6B Yes
Text-to-image, inpainting,
image manipulation
FLUX [103] - 12B Yes
Text-to-image, inpainting,
outpainting, generate variations
InstructPix2Pix [104] - - Yes
Mask-free editing
@@
Figure 20: Images generated by DALL-E 3, SDXL, and Midjourney using the same prompt. (Top row) A cat in a detailed Renaissance cityscape with historical architecture. (Middle row) Highly aesthetic versions of the top row. (Bottom row) Less aesthetic versions of the top row, potentially criticized by art critics.
To compare the models further on their capability to generate art-like images, we asked the three widely used models to generate images that represent Renaissance paintings. In addition to DALL-E 3 [8] and SDXL [101], we include Midjourney in our comparison due to its popularity in text-to-image generation, despite its lack of publicly shared neural network architectures and code. For our comparison, we use prompts to generate images depicting a cat in a cityscape as a Renaissance painting: one highly aesthetic version and another that could potentially be criticized by art critics. Our aim is to see how well these models capture art styles in relation to aesthetics and art criticism. Fig. 20 shows the results. All three models seem to capture the Renaissance style for the cityscape aspect of the image but seem to have difficulty applying this style to the cat. This could be explained by cityscape being a common art genre also present in Renaissance paintings. On the other hand, cats are less common in cityscape or Renaissance paintings but are highly common in natural images as part of these models’ training dataset. This could explain why the cats are often generated in a more modern drawing style and are disproportionately large compared to the buildings in the images. When we compare the highly aesthetic variants to the neutral prompt images, we can see that the highly aesthetic ones are often enhanced in color, brightness or saturation. [108] found that these image features are positively related to human preferences for images generated with a model that aims to enhance the aesthetic value of an image. In a similar line of research, the importance of generating highly aesthetic images has been emphasized by [109], who propose quality-tuning a latent diffusion model [100] for this purpose. In Fig. 20, the images for the art criticism prompts cannot be distinguished easily from the neutral prompts, indicating that the models have a weak notion of art criticism, or that they (most probably correctly) assume that any AI-generated image would be criticized by art critics.
@@
Figure 21: Example images generated by SDXL and Midjourney, showcasing the capability to create (partially or fully) nude images. (Left) Images generated by SDXL. (Right) Images generated by Midjourney. The images in the first and second rows were generated with the prompt “A painting of Snow White in the style of The Birth of Venus by Sandro Botticelli.” The images in the third and fourth rows were generated with the prompt “Batman as a Greek God wearing a loincloth.”